# Towards Domain Adaptive Neural Contextual Bandits

### v1

@@ 4 Method
where $\mathbb{E}^{\mathcal{S}}$ and $\mathbb{E}^{\mathcal{T}}$ denote expectations over the data distributions of $({\bf x},a)$ in the source and target domains, respectively. $g$ is a discriminator that classifies whether ${\bf x}$ is from the source domain or the target domain. $\mathcal{L}_{D}$ is the binary classification accuracy, where labels 00 and $1$ indicate the source and target domains, respectively. In practice, we use the cross-entropy loss as a differentiable surrogate loss. As in [28], solving the minimax optimization $\min\nolimits_{\widehat{\phi}}\max\nolimits_{g}{\mathcal{L}}^{div}$ is equivalent to aligning source- and target-domain data distributions in the latent (encoding) space induced by the encoder $\widehat{\phi}$.
Metrics Acc Per Class $\uparrow$ Acc $\uparrow$
Test Accuracy 0 1 2 3 4 5 6 7 8 9 Average
LinUCB [1] 0.3667$\pm$0.03 0.4045$\pm$0.02 0.5102$\pm$0.03 0.3811$\pm$0.01 0.3157$\pm$0.04 0.3445$\pm$0.01 0.4229$\pm$0.03 0.4595$\pm$0.03 0.3489$\pm$0.02 0.2449$\pm$0.01 0.3808$\pm$0.02
LinUCB-P [1] 0.3645$\pm$0.02 0.6398$\pm$0.04 0.3151$\pm$0.02 0.2709$\pm$0.03 0.2258$\pm$0.01 0.1995$\pm$0.02 0.1752$\pm$0.01 0.3786$\pm$0.04 0.1273$\pm$0.02 0.2992$\pm$0.03 0.3060$\pm$0.02
NLinUCB [11] 0.3781$\pm$0.04 0.3228$\pm$0.03 0.3569$\pm$0.03 0.3381$\pm$0.04 0.3663$\pm$0.04 0.4312$\pm$0.05 0.4766$\pm$0.05 0.4004$\pm$0.05 0.4034$\pm$0.02 0.3613$\pm$0.03 0.3816$\pm$0.04
NLinUCB-P [11] 0.8588$\pm$0.03 0.2224$\pm$0.02 0.3880$\pm$0.03 0.3029$\pm$0.02 0.2978$\pm$0.03 0.0000$\pm$0.01 0.1869$\pm$0.02 0.2243$\pm$0.01 0.2148$\pm$0.01 0.0000$\pm$0.00 0.2706$\pm$0.02
DABand (Ours) 0.6879$\pm$0.10 0.6654$\pm$0.03 0.6056$\pm$0.01 0.5694$\pm$0.05 0.5978$\pm$0.05 0.5428$\pm$0.03 0.6086$\pm$0.02 0.5700$\pm$0.01 0.6261$\pm$0.04 0.5016$\pm$0.02 0.5984$\pm$0.02
Ours vs. NLinUCB +0.3098 +0.3426 +0.2487 +0.2313 +0.2315 +0.1116 +0.1320 +0.1696 +0.2227 +0.1403 +0.2168
Metrics Acc Per Class $\uparrow$ Acc $\uparrow$
LinUCB [1] N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
LinUCB-P [1] 0.0510$\pm$0.02 0.0190$\pm$0.00 0.1360$\pm$0.05 0.0440$\pm$0.02 0.0390$\pm$0.01 0.5580$\pm$0.06 0.0410$\pm$0.02 0.2050$\pm$0.05 0.0850$\pm$0.04 0.0920$\pm$0.13 0.0580$\pm$0.02 0.0460$\pm$0.12 0.1145$\pm$0.06
NLinUCB [11] 0.2870$\pm$0.03 0.0220$\pm$0.02 0.0000$\pm$0.00 0.1250$\pm$0.02 0.0000$\pm$0.00 0.0130$\pm$0.02 0.0470$\pm$0.02 0.0000$\pm$0.00 0.0460$\pm$0.02 0.0050$\pm$0.00 0.6550$\pm$0.04 0.0000$\pm$0.00 0.1001$\pm$0.02
NLinUCB-P [11] 0.0060$\pm$0.01 0.0020$\pm$0.01 0.9222$\pm$0.02 0.0010$\pm$0.00 0.0080$\pm$0.01 0.0990$\pm$0.03 0.0030$\pm$0.00 0.0020$\pm$0.00 0.0180$\pm$0.02 0.0420$\pm$0.01 0.0000$\pm$0.00 0.0150$\pm$0.01 0.0932$\pm$0.01
DABand (Ours) 0.5500$\pm$0.02 0.3560$\pm$0.04 0.4500$\pm$0.02 0.3850$\pm$0.01 0.5590$\pm$0.04 0.2650$\pm$0.03 0.6480$\pm$0.03 0.2410$\pm$0.03 0.4880$\pm$0.04 0.4060$\pm$0.02 0.7680$\pm$0.05 0.1970$\pm$0.02 0.4428$\pm$0.03
Ours vs. NLinUCB +0.2630 +0.3340 +0.4500 +0.2600 +0.5590 +0.2520 +0.6010 +0.2410 +0.4420 +0.4010 +0.1130 +0.1970 +0.3427
Metrics Acc Per Class $\uparrow$ Acc $\uparrow$
LinUCB [1] N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
LinUCB-P [1] 0.1430$\pm$0.02 0.2600$\pm$0.03 0.0810$\pm$0.01 0.0240$\pm$0.01 0.1070$\pm$0.02 0.0570$\pm$0.01 0.1060$\pm$0.01 0.0390$\pm$0.01 0.0960$\pm$0.02 0.1260$\pm$0.03 0.1039$\pm$0.02
NLinUCB [11] 0.0350$\pm$0.00 0.1210$\pm$0.02 0.0510$\pm$0.01 0.0870$\pm$0.02 0.0270$\pm$0.00 0.0240$\pm$0.01 0.2720$\pm$0.04 0.0090$\pm$0.00 0.1730$\pm$0.02 0.0309$\pm$0.01 0.1108$\pm$0.02
NLinUCB-P [11] 0.0000$\pm$0.00 0.0000$\pm$0.00 0.0000$\pm$0.00 0.2990$\pm$0.03 0.7000$\pm$0.02 0.0000$\pm$0.00 0.0000$\pm$0.00 0.0000$\pm$0.00 0.0000$\pm$0.00 0.1160$\pm$0.01 0.1115$\pm$0.01
DABand (Ours) 0.5470$\pm$0.03 0.6410$\pm$0.03 0.5840$\pm$0.03 0.2140$\pm$0.01 0.0460$\pm$0.01 0.7570$\pm$0.04 0.5680$\pm$0.04 0.1340$\pm$0.02 0.1110$\pm$0.02 0.3120$\pm$0.02 0.3914$\pm$0.03
Ours vs. NLinUCB +0.5120 +0.5200 +0.5330 +0.1270 +0.0190 +0.7330 +0.2960 +0.1250 -0.0620 +0.2811 +0.2806
Predicted Reward. The predicted reward term in Eqn. 5, i.e., $\sum\nolimits_{i=1}^{N}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}$, can be translated to the loss term $\sum\nolimits_{i=1}^{N}\mathcal{L}_{i}^{reg}$, where for each round $i$ we have
$\mathcal{L}_{i}^{reg}=\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}\rangle\right|.$ (8)
@@
Constant Term. In Eqn. 5, $\psi+C$ is the constant term, where $\psi$ is defined in Definition 3.7, and $C$ is a positive constant (see the Appendix for more details).
@@ 4.2 Putting Everything Together: DABand Training Algorithm
${\mathcal{L}}^{DABand}=\sum\nolimits_{i=1}^{N}(\mathcal{L}_{i}^{\mathit{R}_{\mathcal{S}}}+\mathcal{L}_{i}^{\epsilon_{\mathcal{S}}(h)}+\mathcal{L}_{i}^{reg})+\lambda\cdot\mathcal{L}_{i}^{div}.$ (10)
Note that we need to alternate between updating $\widehat{\theta}$ using LinUCB [1] and updating the encoder $\widehat{\phi}$ with the discriminator $g$ in Eqn. 9.
Formally, our algorithm operates as follows: In each iteration, we access the context in the original feature space (i.e., $\mathbb{R}^{d}$) for each action from both source and target domains, denoted as $\left\{x_{i,a}^{\mathcal{S}}\right\}_{a\in[K]}$ and $\left\{x_{i,a}^{\mathcal{T}}\right\}_{a\in[K]}$ respectively. Subsequently, we compute the latent representation using the encoder $\phi(\cdot)$ and select the action $\widehat{a}_{i}$ for iteration $i$ based on the LinUCB selection rule. The bandit parameters (i.e., $\theta,A,b$) are updated in each iteration. The weights in $\phi$ are updated every episode of length $H$ (i.e., a batch with $H$ past iterations) using our objective function in Eqn. 9, following the same rule as in Neural-LinUCB [11]. Alg. 1 summarizes our algorithm.
@@ 5 Experiments
Datasets. To demonstrate the effectiveness of our DABand, We evaluate our methods in terms of prediction accuracy and zero-shot target regret on three datasets, i.e., DIGIT [37], VisDA17 [38], and S2RDA49 [39]. See details for each dataset in Appendix B.
Baselines. We compare our DABand with both classic and state-of-the-art contextual bandit algorithms, including LinUCB [1], LinUCB with principle component analysis (PCA) pre-processing (i.e., LinUCB-P), Neural-LinUCB [11] (NLinUCB), and Neural-LinUCB variant that incorporates PCA (i.e., NLinUCB-P). Details for each baselines and discussion are in Appendix C. Note that domain adaptation baselines are not applicable to our setting because it only works in offline settings and assumes complete observability of labels in the source domain (see Appendix C for details).
Zero-Shot Target Regret (Accuracy). We evaluate different methods on the zero-shot target regret setting in Definition 3.5, where all methods can only collect feedback from the source domain, but not the target domain. In this setting, accuracy is equal to $1$ minus the average target regret, i.e., $ACC=1-\tfrac{1}{N}R_{{\mathcal{T}}}$, where $R_{{\mathcal{T}}}$ is defined in Eqn. 2 and Definition 3.5. Therefore higher accuracy indicates lower target regret and better performance.
@@ 6 Conclusion
In this paper, we introduce a novel domain adaptive neural contextual bandit algorithm, DABand, which adeptly combines effective exploration with representation alignment, utilizing unlabeled data from both source and target domains. Our theoretical analysis demonstrates that DABand can attain a sub-linear regret bound within the target domain. This marks the first regret analysis for domain adaptation in contextual bandit problems incorporating deep representation, shallow exploration, and adversarial alignment. We show that all these elements are instrumental in the domain adaptive bandit setting on real-world datasets. Moving forward, interesting future research could be to uncover more innovative techniques for aligning the source and target domains (mentioned in Appendix E), particularly within the constraints of: (1) bandit settings, as opposed to classification settings, and (2) the high-dimensional and dense nature of the target domain, contrasted with the sparse and simplistic nature of the source domain.
@@ A.1 Proofs for Lemmas
By definition of $\mathcal{H}\Delta\mathcal{H}$, we have
$\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})$ $=2\enspace\sup_{h,h^{\prime}\in\mathcal{H}}|\mathbb{E}_{{\bf x}\sim\mathcal{S}}[|h({\bf x})-h^{\prime}({\bf x})|]-\mathbb{E}_{{\bf x}\sim\mathcal{T}}[|h({\bf x})-h^{\prime}({\bf x})|]|$
@@
$=\dfrac{2}{N}\enspace\sup_{h,h^{\prime}\in\mathcal{H}}\Big{|}\epsilon_{\mathcal{S}}(h,h^{\prime})-\epsilon_{\mathcal{T}}(h,h^{\prime})\Big{|}$
@@
$=\sum\nolimits_{i=1}^{N}\bigg{|}h_{1}(x^{\mathcal{D}})-h_{2}(x^{\mathcal{D}})+h_{3}(x^{\mathcal{D}})-h_{3}(x^{\mathcal{D}})\bigg{|}$
@@
$\leq\epsilon_{\mathcal{T}}(h^{*})+\epsilon_{\mathcal{S}}(h,h^{*})+\dfrac{N}{2}\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})$

### v2

@@ 4 Method
where $\mathbb{E}^{\mathcal{S}}$ and $\mathbb{E}^{\mathcal{T}}$ denote expectations over the data distributions of $({\bf x},a)$ in the source and target domains, respectively. $g$ is a discriminator that classifies whether ${\bf x}$ is from the source domain or the target domain. $\mathcal{L}_{D}$ is the binary classification accuracy, where labels 00 and $1$ indicate the source and target domains, respectively. In practice, we use the cross-entropy loss as a differentiable surrogate loss. As in Ganin et al. (2016b), solving the minimax optimization $\min\nolimits_{\widehat{\phi}}\max\nolimits_{g}{\mathcal{L}}^{div}$ is equivalent to aligning source- and target-domain data distributions in the latent (encoding) space induced by the encoder $\widehat{\phi}$.
Predicted Reward. The predicted reward term in Eqn. (5), i.e., $\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right)$, can be translated to the loss term $\sum\nolimits_{i=1}^{N}\mathcal{L}_{i}^{reg}$, where for each round $i$ we have
$\mathcal{L}_{i}^{reg}=\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right).$ (8)
@@
Constant Term. In Eqn. (5), $\psi+C$ is the constant term, where $\psi$ is defined in Definition 3.7, and $C$ contains some small constants which can be ignored (see the Appendix for more details).
@@ 4.2 Putting Everything Together: DABand Training Algorithm
${\mathcal{L}}^{DABand}=\sum\nolimits_{i=1}^{N}(\mathcal{L}_{i}^{\mathit{R}_{\mathcal{S}}}+2\cdot\mathcal{L}_{i}^{\epsilon_{\mathcal{S}}(h)}+\mathcal{L}_{i}^{reg})+\lambda\cdot\mathcal{L}_{i}^{div}.$ (10)
Note that we need to alternate between updating $\widehat{\theta}$ using LinUCB Li et al. (2010) and updating the encoder $\widehat{\phi}$ with the discriminator $g$ in Eqn. (9).
Formally, our method, described in Alg. 1, operates as follows: In each iteration, we access the context in the original feature space (i.e., $\mathbb{R}^{d}$) for each action from both source and target domains, denoted as $\left\{x_{i,a}^{\mathcal{S}}\right\}_{a\in[K]}$ and $\left\{x_{i,a}^{\mathcal{T}}\right\}_{a\in[K]}$ respectively. Subsequently, we compute the latent representation using the encoder $\phi(\cdot)$ and select the action $\widehat{a}_{i}$ for iteration $i$ based on the LinUCB selection rule. The bandit parameters (i.e., $\theta,A,b$) are updated in each iteration. The weights in $\phi$ are updated every episode of length $H$ (i.e., a batch with $H$ past iterations) using our objective function in Eqn. (9), following the same rule as in Neural-LinUCB (Xu et al., 2020).
@@ 5 Experiments
Datasets. To demonstrate the effectiveness of our DABand, We evaluate our methods in terms of prediction accuracy and zero-shot target regret on three datasets, i.e., DIGIT (Ganin et al., 2016a), VisDA17 (Peng et al., 2017), and S2RDA49 (Tang & Jia, 2023). See details for each dataset in Appendix B.
Baselines. We compare our DABand with both classic and state-of-the-art contextual bandit algorithms, including LinUCB (Li et al., 2010), LinUCB with principle component analysis (PCA) pre-processing (i.e., LinUCB-P), Neural-LinUCB (Xu et al., 2020) (NLinUCB), and Neural-LinUCB variant that incorporates PCA (i.e., NLinUCB-P). Details for each baselines and discussion are in Appendix C. Note that domain adaptation baselines are not applicable to our setting because it only works in offline settings and assumes complete observability of labels in the source domain (see Appendix C for details).
Metrics Acc Per Class $\uparrow$ Acc $\uparrow$
Test Accuracy 0 1 2 3 4 5 6 7 8 9 Average
LinUCB (Li et al., 2010) 0.3667$\pm$0.03 0.4045$\pm$0.02 0.5102$\pm$0.03 0.3811$\pm$0.01 0.3157$\pm$0.04 0.3445$\pm$0.01 0.4229$\pm$0.03 0.4595$\pm$0.03 0.3489$\pm$0.02 0.2449$\pm$0.01 0.3808$\pm$0.02
LinUCB-P (Li et al., 2010) 0.3645$\pm$0.02 0.6398$\pm$0.04 0.3151$\pm$0.02 0.2709$\pm$0.03 0.2258$\pm$0.01 0.1995$\pm$0.02 0.1752$\pm$0.01 0.3786$\pm$0.04 0.1273$\pm$0.02 0.2992$\pm$0.03 0.3060$\pm$0.02
NLinUCB (Xu et al., 2020) 0.3781$\pm$0.04 0.3228$\pm$0.03 0.3569$\pm$0.03 0.3381$\pm$0.04 0.3663$\pm$0.04 0.4312$\pm$0.05 0.4766$\pm$0.05 0.4004$\pm$0.05 0.4034$\pm$0.02 0.3613$\pm$0.03 0.3816$\pm$0.04
NLinUCB-P (Xu et al., 2020) 0.8588$\pm$0.03 0.2224$\pm$0.02 0.3880$\pm$0.03 0.3029$\pm$0.02 0.2978$\pm$0.03 0.0000$\pm$0.01 0.1869$\pm$0.02 0.2243$\pm$0.01 0.2148$\pm$0.01 0.0000$\pm$0.00 0.2706$\pm$0.02
DABand (Ours) 0.6891$\pm$0.10 0.6662$\pm$0.03 0.6188$\pm$0.01 0.5703$\pm$0.05 0.6008$\pm$0.05 0.5534$\pm$0.03 0.6010$\pm$0.02 0.5709$\pm$0.01 0.6266$\pm$0.04 0.5023$\pm$0.02 0.6002$\pm$0.02
Ours vs. NLinUCB +0.3110 +0.3434 +0.2619 +0.2322 +0.2345 +0.1222 +0.1244 +0.1705 +0.2232 +0.1410 +0.2186
Metrics Acc Per Class $\uparrow$ Acc $\uparrow$
LinUCB (Li et al., 2010) N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
LinUCB-P (Li et al., 2010) 0.0510$\pm$0.02 0.0190$\pm$0.00 0.1360$\pm$0.05 0.0440$\pm$0.02 0.0390$\pm$0.01 0.5580$\pm$0.06 0.0410$\pm$0.02 0.2050$\pm$0.05 0.0850$\pm$0.04 0.0920$\pm$0.13 0.0580$\pm$0.02 0.0460$\pm$0.12 0.1145$\pm$0.06
NLinUCB (Xu et al., 2020) 0.2870$\pm$0.03 0.0220$\pm$0.02 0.0000$\pm$0.00 0.1250$\pm$0.02 0.0000$\pm$0.00 0.0130$\pm$0.02 0.0470$\pm$0.02 0.0000$\pm$0.00 0.0460$\pm$0.02 0.0050$\pm$0.00 0.6550$\pm$0.04 0.0000$\pm$0.00 0.1001$\pm$0.02
NLinUCB-P (Xu et al., 2020) 0.0060$\pm$0.01 0.0020$\pm$0.01 0.9222$\pm$0.02 0.0010$\pm$0.00 0.0080$\pm$0.01 0.0990$\pm$0.03 0.0030$\pm$0.00 0.0020$\pm$0.00 0.0180$\pm$0.02 0.0420$\pm$0.01 0.0000$\pm$0.00 0.0150$\pm$0.01 0.0932$\pm$0.01
DABand (Ours) 0.5504$\pm$0.02 0.3562$\pm$0.04 0.4462$\pm$0.02 0.3844$\pm$0.01 0.5582$\pm$0.04 0.2632$\pm$0.03 0.6474$\pm$0.03 0.2396$\pm$0.03 0.4882$\pm$0.04 0.4062$\pm$0.02 0.7685$\pm$0.05 0.1968$\pm$0.02 0.4644$\pm$0.03
Ours vs. NLinUCB +0.2634 +0.3342 +0.4462 +0.2594 +0.5582 +0.2502 +0.6004 +0.2396 +0.4422 +0.4012 +0.1135 +0.1968 +0.3643
Zero-Shot Target Regret (Accuracy). We evaluate different methods on the zero-shot target regret setting in Definition 3.5, where all methods can only collect feedback from the source domain, but not the target domain. In this setting, accuracy is equal to $1$ minus the average target regret, i.e., $ACC=1-\tfrac{1}{N}R_{{\mathcal{T}}}$, where $R_{{\mathcal{T}}}$ is defined in Eqn. (2) and Definition 3.5. Therefore higher accuracy indicates lower target regret and better performance.
@@
Metrics Acc Per Class $\uparrow$ Acc $\uparrow$
LinUCB (Li et al., 2010) N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
LinUCB-P (Li et al., 2010) 0.1430$\pm$0.02 0.2600$\pm$0.03 0.0810$\pm$0.01 0.0240$\pm$0.01 0.1070$\pm$0.02 0.0570$\pm$0.01 0.1060$\pm$0.01 0.0390$\pm$0.01 0.0960$\pm$0.02 0.1260$\pm$0.03 0.1039$\pm$0.02
NLinUCB (Xu et al., 2020) 0.0350$\pm$0.00 0.1210$\pm$0.02 0.0510$\pm$0.01 0.0870$\pm$0.02 0.0270$\pm$0.00 0.0240$\pm$0.01 0.2720$\pm$0.04 0.0090$\pm$0.00 0.1730$\pm$0.02 0.0309$\pm$0.01 0.1108$\pm$0.02
NLinUCB-P (Xu et al., 2020) 0.0000$\pm$0.00 0.0000$\pm$0.00 0.0000$\pm$0.00 0.2990$\pm$0.03 0.7000$\pm$0.02 0.0000$\pm$0.00 0.0000$\pm$0.00 0.0000$\pm$0.00 0.0000$\pm$0.00 0.1160$\pm$0.01 0.1115$\pm$0.01
DABand (Ours) 0.5501$\pm$0.03 0.6418$\pm$0.03 0.5844$\pm$0.03 0.2346$\pm$0.01 0.0410$\pm$0.01 0.7580$\pm$0.04 0.5689$\pm$0.04 0.1345$\pm$0.02 0.1118$\pm$0.02 0.3126$\pm$0.02 0.3923$\pm$0.03
Ours vs. NLinUCB +0.5151 +0.5208 +0.5334 +0.1476 +0.0140 +0.7340 +0.2969 +0.1255 -0.0612 +0.2817 +0.2815
@@ 6 Conclusion
In this paper, we introduce a novel domain adaptive neural contextual bandit algorithm, DABand, which adeptly combines effective exploration with representation alignment, utilizing unlabeled data from both source and target domains. Our theoretical analysis demonstrates that DABand can attain a sub-linear regret bound within the target domain. This marks the first regret analysis for domain adaptation in contextual bandit problems incorporating deep representation, shallow exploration, and adversarial alignment. We show that all these elements are instrumental in the domain adaptive bandit setting on real-world datasets. Moving forward, interesting future research could be to uncover more innovative techniques for aligning the source and target domains (mentioned in Appendix F), particularly within the constraints of: (1) bandit settings, as opposed to classification settings, and (2) the high-dimensional and dense nature of the target domain, contrasted with the sparse and simplistic nature of the source domain.
@@ A.1 Proofs for Lemmas
By definition of $\mathcal{H}\Delta\mathcal{H}$ (as defined in Def. 3.6), we have
[…]
@@
[…]
@@
[…]
@@
[…]